OpenStudy (anonymous):

An investor invested in a total of $3200 in two mutual funds. One fund is 5% profit and other is 2% profit. If the total profit was $130, then how much was invested in the mutual fund? 5% =? and 3% =?

OpenStudy (anonymous):

so you would write system of equation

OpenStudy (anonymous):

Okay how would I go about doing this

OpenStudy (anonymous):

wait is it 2% or 3%?

OpenStudy (anonymous):

3%

OpenStudy (anonymous):

5% and 3%

OpenStudy (anonymous):

let define a= amount invested in 3% b= amount invested in 5%

OpenStudy (anonymous):

a+b=3200 make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now another equation .03 * a+.05 b=130

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

.03a+.0b=130?

OpenStudy (anonymous):

.05

OpenStudy (anonymous):

3 percent is .03 in decimal if you multiply by amount invested , you get profit

OpenStudy (anonymous):

.03 x 130?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

130 is total invest or 3200

OpenStudy (anonymous):

3200 is total investment

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

a + b = 3200 .03a+.05b=130 now you have to solve for a and b

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

do it

OpenStudy (anonymous):

5x+3(3200-x)=130?

OpenStudy (anonymous):

a+b=3200 a=3200-b

OpenStudy (anonymous):

sub that into second equation

OpenStudy (anonymous):

.03a+.05b=130 .03(3200-b)+.05b=130

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

I am so illiterate with these types of math problems